# UDPNet: Unleashing Depth-based Priors for Robust Image Dehazing

### v1

@@ 2.2 Depth Priors for Image Dehazing
End-to-end dehazing networks relying solely on RGB features have made significant progress but often struggle with ambiguous scene structures and insufficient depth information constraints. In contrast, dehazing approaches Yang and Zhang (2022); Yang et al (2022); Ye et al (2025) incorporating depth maps offer distinct advantages by leveraging the structure priors embedded in data. The intrinsic physical correlation further highlights the potential. Lee et al. Lee et al (2020) propose a CNN-based simultaneous dehazing and depth estimation network with a shared encoder and multiple decoders. Similarly, Yang et al. Yang and Zhang (2022) and Cheng et al. Cheng and Zhao (2021) propose the depth-aware methods to fuse the depth features to the dehazing network. Due to the absence of accurate depth maps during inference and the possible residual haze remaining, Wang et al. Wang et al (2024a) propose a continuous self-prompting inference strategy, which iteratively refines the dehazing results toward clearer outputs. Furthermore, PromptHaze Ye et al (2025) introduces a dehazing strength factor to achieve a controllable step-by-step dehazing process, thereby alleviating the impact of cumulative errors. Recently, Zhang et al. Zhang et al (2024) integrate lightweight depth estimation model and image dehazing to form a dual-task collaborative optimization. Some novel methods Yang et al (2022); Wu et al (2023); Liu et al (2025); Ye et al (2025) perform online synthesis of realistic hazy images based on depth maps and physical models, facilitating supervised training for real-world dehazing. These innovations inspire us to develop an effective and general dehazing framework that leverages depth priors as guidance. Moreover, our work proposes a general depth map fusion scheme, which is validated through experiments on previous algorithms Cui et al (2023b, 2024a, 2025a); Potlapalli et al (2023); Cui et al (2025b). In the following sections, we will introduce the detailed design and implementation of our methods.
Figure 3: Cross-dataset statistical visualization of residual haze (i.e., haze density) across depth ranges highlighting scene-dependent variations and depth–haze correlations in indoor, outdoor, nighttime, remote sensing, and real-world scenarios.
2.3 DepthAnything vs. Other Depth Estimation Models
@@
DepthAnything Yang et al (2024a) and its successor DepthAnything V2 Yang et al (2024b) represent a significant shift in monocular depth estimation through a precise synthetic and large‑scale pseudo‑labeled paradigm. Unlike MiDaSv3.1 Birkl et al (2023) or AdaBins Bhat et al (2021), which are exclusively trained on manually annotated datasets (e.g., NYU‑D Silberman et al (2012), KITTI Geiger et al (2013)), DepthAnything first distilled a high-capacity DINOv2-Giant Oquab et al (2023) teacher model on five high‐fidelity synthetic datasets, and then generated pseudo‐depth for 62 million in‑the‑wild real images, effectively bridging the domain gap between synthetic and real-world data distributions. Finally, compact student models (ViT-S/B/L/G variants) were trained on these high-quality pseudo-labels. This methodology leads to superior performance in handling thin structures and transparent objects, and significantly improves generalization across complex and challenging scenes. The global depth scene information, coupled with the refined sharpness of depth discontinuities, effectively serves as an auxiliary to image dehazing tasks. Our framework ensures a seamless integration of DepthAnything V2 priors with multi-scale dehazing networks.
@@
Figure 4: Comparative analysis of the impact of depth map assistance on various image restoration tasks (i.e., dehazing, deraining, denoising, deblurring, and low-light enhancement). Our experiments show that incorporating depth maps yields substantial gains in dehazing and low-light enhancement, demonstrating their potential for both daytime and nighttime dehazing. Moreover, the overall performance achieves state-of-the-art results. Depth maps are generated using DepthAnything-small Yang et al (2024a), and our dehazing network, PromptIR Potlapalli et al (2023), takes both RGB and depth inputs rather than relying solely on RGB features.
3 Preliminary Analysis
@@
It suggests that deeper regions should appear more heavily degraded. However, in practice this monotonic trend does not always hold across different scenes. To better understand this discrepancy, we analyze the relationship between depth and residual haze, where residual haze is defined as the absolute difference between the hazy input and its corresponding clear image. Specifically, for each dataset we compute a residual haze map, partition the normalized depth into three equal bins (Near, Mid, Far), and calculate the average residual haze intensity within each bin. Fig. 3 presents a cross-dataset statistical analysis. Interestingly, the trends vary significantly across scene categories. Outdoor datasets exhibit stronger haze density in distant regions, which is consistent with physical scattering intuition, whereas indoor datasets show stronger residuals at near distances. Nighttime datasets present different residuals, reflecting the compounded degradations from haze and illumination. Remote sensing scenes reveal consistently strong haze density across all depths, while real-world datasets demonstrate a mixed trend. These observations highlight that the relationship between depth and haze density is scene-dependent, emphasizing the need to incorporate depth priors for adaptive dehazing, rather than relying on handcrafted priors or rigid physical formula-based constraints.
@@
Meanwhile, we further argue that depth estimation models inherently have the capacity to jointly handle dehazing and low-light enhancement, and this capacity can be further enhanced through elaborate design. In all-in-one image restoration, a single network is expected to address diverse degradations and adaptively decide how to restore the input. Depth information provides valuable guidance, particularly for dehazing and low-light scenarios, by separating foreground from background and enabling more accurate restoration across varying depths. This improves robustness in complex real-world conditions. To validate the effectiveness of our proposed method, we perform preliminary experiments for verification. We experiment with a dedicated all-in-one image restoration network Potlapalli et al (2023) evaluating on the five-degradation scenarios Jiang et al (2025) (modifying the original RGB input by concatenating it with depth maps extracted by DepthAnything-small Yang et al (2024a)). The results are presented in Fig. 4. It can be seen that the average PSNR achieves state-of-the-art performance, with particularly significant improvements on the dehazing and low-light enhancement tasks. This indicates that the depth maps extracted by DepthAnything inherently contribute to both daytime and nighttime image dehazing. Additionally, the depth-based approach shows significant potential for enhancing performance in other image restoration tasks, deserving further exploration.
@@
Figure 5: The architecture of our proposed UDPNet (a) is inspired by the symmetric hierarchical encoder-decoder design Cui et al (2023b). UDPNet mainly comprises two organically integrated sub-modules, i.e., DGAM (b) and DPFM (d). DGAM focuses on channel-level fusion between depth-based priors and the input image. DPFM incorporates a multi-scale, multi-head cross-attention mechanism with overlapping windows and dual attention pathways at the spatial level. This design enables UDPNet to act as a general paradigm that leverages depth information to boost existing dehazing models, achieving a trade-off between efficiency and effectiveness.
4 Methodology
@@
In this section, we introduce our proposed depth-based framework, as illustrated in Fig. 4. We first present an overall pipeline of our UDPNet. We then delineate details of our designs: the Depth-Guided Attention Module (detailed in Sec. 4.2.1) and the Depth Prior Fusion Module (detailed in Sec. 4.2.2). The loss functions are introduced in the final part.
4.1 Overall Architecture
@@
In image dehazing, depth information has proven to be a valuable cue for recovering structural details and improving robustness, especially in complex hazy environments. However, traditional methods often struggle to fully exploit depth priors due to challenges in accurately estimating depth and integrating it effectively with image features. Motivated by this, we aim to integrate depth-based priors into a U-shaped convolutional neural network (CNN) backbone for haze removal. Our approach leverages a pretrained depth estimation model, DepthAnything V2, to generate robust depth maps, which are then used to guide the feature extraction process, enhancing the network’s ability to capture geometric and structural information. By incorporating these priors into an encoder-decoder architecture, we ensure that both high-level and low-level features are effectively utilized, with particular attention to how depth information can aid in early-stage feature extraction.
@@
Each stage of the encoder includes a Depth Prior Fusion Module (DPFM), designed to better integrate depth priors with the visual features. To ensure effective integration, this module is applied in each encoder scale, as depth map information is most beneficial for capturing geometric cues and structural details during the early stages of feature extraction. Introducing the module in the decoder would inject low-frequency information from the depth map into the high-frequency image restoration process, potentially hindering fine-grained detail restoration.
Next, the depth-enhanced features are passed through a three-scale decoder, progressively restoring high-resolution features. During training, a multi-output strategy is employed, where low-resolution clean images are predicted using $3\times 3$ convolutions and image-level skip connections. For brevity, Fig. 5(a) only shows the top-level image skip connection. Although our approach can be applied to various dehazing baselines, we demonstrate the framework on FSNet Cui et al (2023b) for clarity.
4.2 Depth Priors Fusion
Effective image dehazing requires not only removal of haze but also faithful recovery of underlying scene structures, spatial consistency, and depth-aware contrasts. While RGB images provide rich color information, they often lack sufficient cues to resolve structural ambiguities, especially in complex haze conditions. Depth priors, capturing geometric layouts and relative distances within a scene, offer complementary information that can guide the network toward more accurate restoration. However, existing methods either rely on simple concatenation of depth maps with image features or apply simple integration strategies, which fail to fully exploit the rich structural knowledge embedded in depth cues. Motivated by this, we aim to systematically incorporate depth priors from a large-scale pretrained model into the dehazing pipeline, enhancing both fidelity and generalization across diverse haze scenarios.
@@ 4.2 Depth Priors Fusion
4.2.1 Depth-Guided Attention Module
@@
The refined depth map is concatenated with the RGB image and fed into the subsequent module to extract depth-aware features. We first employ a $3\times 3$ convolution followed by instance normalization and GELU activation to process the concatenated RGB-D input and produce intermediate feature representations. To further enhance feature discrimination, a depth-guided attention mechanism is applied. This module captures global context through adaptive average pooling, generates channel-wise attention weights via two fully connected layers with a GELU nonlinearity, and applies a sigmoid activation. The resulting attention weights are then used to reweight the input features, thereby emphasizing informative channels and suppressing less relevant ones. Formally, the process can be expressed as
@@
where $\cdot$$\cdot$] is a concatenation operator, $\odot$ denotes element-wise multiplication, GAP is global average pooling, $\text{FC}_{1}$ and $\text{FC}_{2}$ are fully connected layers, and $\sigma$ is the sigmoid activation. By integrating the refined depth map and RGB image through this DGAM, as depicted in Fig. 5(b), the network obtains depth-aware shallow features that benefit subsequent restoration stages. The objective experimental results validate that the proposed DGAM is more effective than a simple concatenation strategy. As detailed in Section 5.4, employing DGAM yields a performance gain on the Haze4K dataset compared to naive feature stacking, demonstrating the advantage of adaptive, depth-guided channel attention.
4.2.2 Depth Prior Fusion Module
The representations encoded in shallow and deep features differ substantially, as they correspond to entirely different receptive fields. This disparity makes it essential to unleash depth priors across multiple feature scales. Inspired by SwinIR Liang et al (2021), HAT Chen et al (2023) and Restormer Zamir et al (2022), we propose a depth prior fusion module (DPFM) to inject estimated depth priors into the multi-scale restoration backbone via a dual cross-attention mechanism defined on overlapping local windows, as shown in Fig. 5(d). Given an intermediate feature map $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$ and a single-channel depth prior $\mathbf{D}\in\mathbb{R}^{H\times W\times 1}$, the depth map is first projected by shallow feature extract to produce depth map features $\mathbf{F}_{D}\in\mathbb{R}^{H\times W\times C}$. Both $\mathbf{X}$ and $\mathbf{F}_{D}$ are partitioned into local windows of size $M\times M$ with overlap ratio $r$ (so the effective overlap-window side length is $M_{\mathrm{ov}}=(1+r)\times M$). Zero-padding is first applied to make $H$ and $W$ divisible by $M$. Queries are then partitioned into non-overlapping windows of size $M\times M$, while keys and values are unfolded from larger overlapping windows of size $M_{\mathrm{ov}}\times M_{\mathrm{ov}}$. This overlap allows each query window to incorporate neighboring context at a controllable cost, thereby propagating geometric cues across adjacent regions while keeping the computation tractable.
@@ 4.2.2 Depth Prior Fusion Module
In practice, a learnable relative-position bias is introduced to encode spatial priors. Compared to global cross-attention with quadratic cost $\mathcal{O}((HW)^{2})$, the window overlapping design yields complexity $\mathcal{O}(HW\cdot M\cdot M_{\mathrm{ov}})$, since each $M^{2}$ query interacts with $M_{\mathrm{ov}}^{2}$ key/value tokens. Although slightly higher than standard non-overlapping window cross attention ($\mathcal{O}(HW\cdot M^{2})$), the complexity remains linear in the number of pixels $HW$ and quadratic only in the local window sizes, thus available for high-resolution image restoration. Furthermore, we insert DPFM blocks at multiple encoder stages, allowing depth-aware priors to be progressively injected into the backbone. This hierarchical design enhances geometric consistency and structural fidelity in the restored images, fusing the multi-scale depth information at negligible extra computational cost. The ablation results in Section 5.4 indicate that inserting the DPFM within the encoder stages is more beneficial than placing it in the decoder. This is attributed to the fact that depth priors provide strong structural and geometric cues that are most effective when integrated during the early stages of feature extraction, providing better guidance for subsequent hierarchical processing. Furthermore, the spatial cross-attention mechanism in DPFM proves particularly effective, contributing significantly to the overall state-of-the-art performance achieved on Haze4K.
4.3 Loss Functions
Following the Cui et al. Cui et al (2023b, 2024a), we employ multi-scale dual-domain loss to train our UDPNet. The spatial and spectral $\mathcal{L}_{\text{1}}$ loss are given by:
$\mathcal{L}_{\text{spatial}}=\sum_{i=1}^{3}\frac{1}{P_{i}}\|\hat{\mathbf{I}}_{i}-\mathbf{Y}_{i}\|_{1},$ (9)
$\mathcal{L}_{\text{frequency}}=\sum_{i=1}^{3}\frac{1}{S_{i}}\left\|\left[\mathcal{R}(\hat{\mathbf{I}}_{i}),\mathcal{I}(\hat{\mathbf{I}}_{i})\right]-\left[\mathcal{R}(\mathbf{Y}_{i}),\mathcal{I}(\mathbf{Y}_{i})\right]\right\|_{1},$ (10)

### v2

@@ II-B Depth Priors for Image Dehazing
While end-to-end dehazing networks relying solely on RGB features have advanced significantly, they often struggle with ambiguous scene structures due to insufficient depth constraints. In contrast, approaches incorporating depth maps [yang2022depth, yang2022self, ye2025prompthaze] provide distinct advantages by leveraging embedded structural priors. The intrinsic physical correlation further highlights the potential. Lee et al. [lee2020cnn] propose a CNN-based simultaneous dehazing and depth estimation network with a shared encoder and multiple decoders. Similarly, Yang et al. [yang2022depth] and Cheng et al. [cheng2021two] propose the depth-aware methods to fuse the depth features to the dehazing network. Due to the absence of accurate depth maps during inference and the possible residual haze remaining, Wang et al. [wang2024selfpromer] propose a continuous self-prompting inference strategy, which iteratively refines the dehazing results toward clearer outputs. Furthermore, PromptHaze [ye2025prompthaze] introduces a dehazing strength factor to achieve a controllable step-by-step dehazing process, thereby alleviating the impact of cumulative errors. Recently, Zhang et al. [zhang2024depth] integrate lightweight depth estimation model and image dehazing to form a dual-task collaborative optimization. Some novel methods [yang2022self, wu2023ridcp, liu2025peie, ye2025prompthaze] perform online synthesis of realistic hazy images based on depth maps and physical models, facilitating supervised training for real-world dehazing. These innovations inspire us to develop an effective and general dehazing framework that leverages depth priors as guidance. Moreover, our work proposes a general depth map fusion scheme, which is validated through experiments on previous algorithms [cui2023image, cui2024revitalizing, cui2025exploring, potlapalli2023promptir, cui2025adair]. In the following sections, we will introduce the detailed design of our methods.
II-C DepthAnything vs. Other Depth Estimation Models
@@
DepthAnything [depthanything] and its successor DepthAnything V2 [depthanything_v2] represent a significant shift in monocular depth estimation through a precise synthetic and large‑scale pseudo‑labeled paradigm. Unlike MiDaSv3.1 [birkl2023midas] or AdaBins [bhat2021adabins], which are exclusively trained on manually annotated datasets (e.g., NYU‑D [silberman2012indoor], KITTI [geiger2013vision]), DepthAnything first distilled a high-capacity DINOv2-Giant [oquab2023dinov2] teacher model on five high‐fidelity synthetic datasets, and then generated pseudo‐depth for 62 million in‑the‑wild real images, effectively bridging the domain gap between synthetic and real-world data distributions. This methodology leads to superior performance in handling thin structures and transparent objects, and significantly improves generalization across complex and challenging scenes. The global depth scene information, coupled with the refined sharpness of depth discontinuities, effectively serves as an auxiliary to image dehazing tasks. Our framework ensures a seamless integration of depth-based priors with dehazing networks.
Figure 3: Cross-dataset statistical visualization of the residual haze (i.e., haze density) across depth ranges highlighting scene-dependent variations and depth–haze correlations in indoor, outdoor, nighttime, remote sensing, and real-world scenarios.
@@
Figure 4: Comparative analysis of the impact of depth map assistance on various image restoration tasks. Our experiments show that incorporating depth maps yields substantial gains in dehazing and low-light enhancement, demonstrating their potential for both daytime and nighttime dehazing. Moreover, the overall performance achieves state-of-the-art results. Depth maps are generated using DepthAnything-small [depthanything], and our dehazing network, PromptIR [potlapalli2023promptir], takes both RGB and depth inputs rather than relying solely on RGB features.
III Preliminary Analysis
@@
This implies that deeper regions should appear more heavily degraded. However, in practice this monotonic trend does not always hold across different scenes. To better understand this discrepancy, we analyze the relationship between depth and residual haze, where residual haze is defined as the absolute difference between the hazy input and its corresponding clear image. Specifically, for each dataset we compute a residual haze map, partition the normalized depth into three equal bins (Near, Mid, Far), and calculate the average residual haze intensity within each bin. Fig. 3 presents a cross-dataset statistical analysis. Interestingly, the trends vary significantly across scene categories. Outdoor datasets exhibit stronger haze density in distant regions, which is consistent with physical scattering intuition, whereas indoor datasets show stronger residuals at near distances. Nighttime datasets present different residuals, reflecting the compounded degradations from haze and illumination. Remote sensing scenes reveal consistently strong haze density across all depths, while real-world datasets demonstrate a mixed trend. These observations highlight that the relationship between depth and haze density is scene-dependent, emphasizing the need to incorporate depth priors for adaptive dehazing, rather than relying on handcrafted priors or rigid physical formula-based constraints.
@@
Meanwhile, we further argue that depth estimation models have the capacity to jointly handle dehazing and low-light enhancement, and this capacity can be further enhanced through elaborate design. In all-in-one image restoration, a single network is expected to address diverse degradations and adaptively decide how to restore the input. Depth information provides valuable guidance, particularly for dehazing and low-light scenarios, by separating foreground from background and enabling more accurate restoration across varying depths. This improves robustness in complex real-world conditions. To validate the effectiveness of our proposed method, we perform preliminary experiments for verification. We experiment with a dedicated all-in-one image restoration network [potlapalli2023promptir] evaluating on the five-degradation scenarios [jiang2024survey] (modifying the original RGB input by concatenating it with depth maps extracted by DepthAnything-small [depthanything]). The results are presented in Fig. 4. It can be seen that the average PSNR achieves state-of-the-art performance, with particularly significant improvements on the dehazing and low-light enhancement tasks. This indicates that the depth maps extracted by DepthAnything inherently contribute to both daytime and nighttime image dehazing. Additionally, the depth-based approach shows significant potential for enhancing performance in other image restoration tasks, deserving further exploration.
@@
Figure 5: The architecture of our proposed UDPNet (a) is inspired by the symmetric hierarchical encoder-decoder design [cui2023image]. UDPNet is a specific instantiation of our UDP (Unleashing Depth-based Priors) paradigm, which is realized by integrated sub-modules, i.e., DGAM (b) and DPFM (d). DGAM focuses on channel-level fusion between depth-based priors and the input image. DPFM incorporates a multi-scale, multi-head cross-attention mechanism with overlapping windows and dual attention pathways at the spatial level.
IV Methodology
@@
In this section, we introduce our proposed depth-based framework, as illustrated in Fig. 5. We first present an overall pipeline of our UDPNet. We then delineate details of our designs: the Depth-Guided Attention Module (detailed in Sec. IV-B1) and the Depth Prior Fusion Module (detailed in Sec. IV-B2). The loss functions are introduced in the final part.
IV-A Overall Architecture
@@
In image dehazing, depth information has proven to be a valuable cue for recovering structural details and improving robustness, especially in complex hazy environments. However, traditional methods often struggle to fully exploit depth priors due to challenges in accurately estimating depth and integrating it effectively with image features. Motivated by this, we aim to integrate depth-based priors into a U-shaped backbone for haze removal. Our approach leverages a pretrained depth estimation model, DepthAnything V2, to generate robust depth maps, which are then used to guide the feature extraction process, enhancing the network’s ability to capture geometric and structural information. By incorporating these priors into a multi-scale architecture, we ensure that both high-level and low-level features are effectively utilized.
@@
Each stage of the encoder includes a Depth Prior Fusion Module (DPFM), designed to better integrate depth priors with the visual features. To ensure effective integration, this module is applied in the encoder, as depth map information is most beneficial for capturing geometric cues and structural details during the early stages of feature extraction. Introducing the module in the decoder would inject low-frequency information from the depth map into the high-frequency image restoration process, potentially hindering fine-grained detail restoration. Next, the depth-enhanced features are passed through a three-scale decoder, progressively restoring high-resolution features. During training, a multi-output strategy is employed, where low-resolution clean images are predicted using $3\times 3$ convolutions and skip connections. For brevity, Fig. 5(a) only shows the top-level image skip connection. Although our approach can be applied to various dehazing baselines, we demonstrate the framework on FSNet [cui2023image] for clarity.
IV-B Depth Priors Fusion
Effective image dehazing requires not only removal of haze but also faithful recovery of underlying scene structures, spatial consistency, and depth-aware contrasts. While RGB images provide rich color information, they often lack sufficient cues to resolve structural ambiguities, especially in complex haze conditions. Depth priors, capturing geometric layouts and relative distances within a scene, offer complementary information that can guide the network toward more accurate restoration. However, existing methods either rely on simple concatenation of depth maps with image features or apply simple integration strategies, which fail to fully exploit the rich structural knowledge embedded in depth cues. Motivated by this, we aim to systematically incorporate depth priors from a large-scale pretrained model into the dehazing pipeline, enhancing both fidelity and generalization across diverse haze.
@@ IV-B Depth Priors Fusion
IV-B1 Depth-Guided Attention Module
@@
The refined depth map is concatenated with the RGB image and fed into the subsequent modules to extract depth-aware features. To further enhance feature discrimination, a depth-guided attention mechanism is applied. We first employ a $3\times 3$ convolution followed by instance normalization and GELU activation to process the concatenated RGB-D input and produce intermediate feature representations. Subsequent modules capture global context through adaptive average pooling, generate channel-wise attention weights via two fully connected layers with a GELU nonlinearity, and apply a sigmoid activation. The resulting attention weights are then used to reweight the input features, thereby emphasizing informative channels and suppressing less relevant ones. Formally, the process can be expressed as
@@
where $\cdot$$\cdot$] is a concatenation operator, $\odot$ denotes element-wise multiplication, $F_{\text{conv}}$ denotes output features of the convolution block, GAP is global average pooling, $\text{FC}_{1}$ and $\text{FC}_{2}$ are fully connected layers, and $\sigma$ is the sigmoid activation. By integrating the refined depth map and RGB image through this DGAM, as depicted in Fig. 5(b), the network obtains depth-aware shallow features that benefit subsequent restoration stages. The objective experimental results validate that the proposed DGAM is more effective than a simple concatenation strategy. As detailed in Section V-D, employing DGAM yields a performance gain on the Haze4K dataset compared to naive feature stacking, demonstrating the advantage of adaptive, depth-guided channel attention.
IV-B2 Depth Prior Fusion Module
The representations encoded in shallow and deep features differ substantially, as they correspond to entirely different receptive fields. This disparity makes it essential to unleash depth priors across multiple feature scales. Inspired by SwinIR [liang2021swinir] and Restormer [zamir2022restormer], we propose a depth prior fusion module (DPFM) to inject estimated depth priors into the multi-scale restoration backbone via a dual cross-attention mechanism defined on overlapping local windows, as shown in Fig. 5(d). Given an intermediate feature map $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$ and a single-channel depth prior $\mathbf{D}\in\mathbb{R}^{H\times W\times 1}$, the depth map is first projected by shallow feature extract to produce depth map features $\mathbf{F}_{D}\in\mathbb{R}^{H\times W\times C}$. Both $\mathbf{X}$ and $\mathbf{F}_{D}$ are partitioned into local windows of size $M\times M$ with overlap ratio $r$ (so the effective overlap-window side length is $M_{\mathrm{ov}}=(1+r)\times M$). Zero-padding is first applied to make $H$ and $W$ divisible by $M$. Queries are then partitioned into non-overlapping windows of size $M\times M$, while keys and values are unfolded from larger overlapping windows of size $M_{\mathrm{ov}}\times M_{\mathrm{ov}}$. This overlap enables each query window to incorporate neighboring context at a controllable cost. As a result, it propagates geometric cues across adjacent regions while maintaining computational tractability.
@@ IV-B2 Depth Prior Fusion Module
In practice, a learnable relative-position bias is introduced to encode spatial priors. Compared to global cross-attention with quadratic cost $\mathcal{O}((HW)^{2})$, the window overlapping design yields complexity $\mathcal{O}(HW\cdot M\cdot M_{\mathrm{ov}})$, since each $M^{2}$ query interacts with $M_{\mathrm{ov}}^{2}$ key/value tokens. Although slightly higher than standard non-overlapping window cross attention ($\mathcal{O}(HW\cdot M^{2})$), the complexity remains linear in the number of pixels $HW$ and quadratic only in the local window sizes, thus available for high-resolution image restoration. Furthermore, we insert DPFM blocks at multiple encoder stages, allowing depth-aware priors to be progressively injected into the backbone. This hierarchical design enhances geometric consistency and structural fidelity in the restored images, fusing the multi-scale depth information at negligible extra computational cost. The ablation results in Section V-D indicate that inserting the DPFM within the encoder stages is more beneficial than placing it in the decoder. This is attributed to the fact that depth priors provide strong structural and geometric cues that are most effective when integrated during the early stages of feature extraction, providing better guidance for subsequent hierarchical processing. Furthermore, the spatial cross-attention mechanism in DPFM proves particularly effective, contributing significantly to the overall state-of-the-art performance achieved on Haze4K.
IV-C Loss Functions